### v1

@@ Abstract
Short Text Classification (STC) is crucial for processing and comprehending the brief but substantial content prevalent on contemporary digital platforms. The STC encounters difficulties in grasping semantic and syntactic intricacies, an issue that is apparent in traditional pre-trained language models. Although Graph Convolutional Networks enhance performance by integrating external knowledge bases, these methods are limited by the quality and extent of the knowledge applied. Recently, the emergence of Large Language Models (LLMs) and Chain-of-Thought (CoT) has significantly improved the performance of complex reasoning tasks. However, some studies have highlighted the limitations of their application in fundamental NLP tasks. Consequently, this study sought to employ CoT to investigate the capabilities of LLMs in STC tasks. This study introduces Quartet Logic: A Four-Step Reasoning (QLFR) framework. This framework primarily incorporates Syntactic and Semantic Enrichment CoT, effectively decomposing the STC task into four distinct steps: (i) essential concept identification, (ii) common-sense knowledge retrieval, (iii) text rewriting, and (iv) classification. This elicits the inherent knowledge and abilities of LLMs to address the challenges in STC. Surprisingly, we found that QLFR can also improve the performance of smaller models. Therefore, we developed a CoT-Driven Multi-task learning (QLFR-CML) method to facilitate the knowledge transfer from LLMs to smaller models. Extensive experimentation across six short-text benchmarks validated the efficacy of the proposed methods. Notably, QLFR achieved state-of-the-art performance on all datasets, with significant improvements, particularly on the Ohsumed and TagMyNews datasets.
@@ 1 Introduction
Short texts are crucial to the contemporary flow of information, particularly given the rapid growth of the Internet[1]. They are essential on major social media platforms, including Twitter, TikTok, Instagram, and Weibo, where they facilitate social interaction and are integral to daily activities. As a vital task for intelligent empowerment and the application of short texts, short-text classification (STC) is essential for applications such as news categorization[2], question answering (QA)[3], and sentiment analysis[4]. Traditional pre-trained language models (PLMs) struggle with the semantic and syntactic intricacies of STC[1, 5]. As shown in Fig 1, categorizing a news title such as ‘Del Potro says make French Open’ into the ‘sport’ category can be challenging, especially without recognizing ‘Del Potro’ as a professional tennis player. The absence of a clear subject and predicate impedes the model’s understanding and learning capacity. To address the challenges posed in STC tasks, Graph Convolutional Networks (GCNs) have shown a certain degree of progress by using an additional knowledge base and redefining STC as a node classification issue, which mitigates the problem of limited training data. However, the effectiveness of GCNs is limited by the quality and extent of their knowledge.
@@
Figure 1: This diagram compares two approaches to the STC task. Due to misinterpretation, the traditional approach erroneously classifies the input ‘Del Potro says make French Open’ as ‘world’. Conversely, our method employs a sequential analytical process that correctly identifies ‘Del Potro’ as a tennis player, recognizes the ‘French Open’ as a tennis tournament, and detects the absence of a grammatical object in the sentence, resulting in accurate categorization under ‘sport’.
@@
The NLP landscape has been revolutionized by Large Language Models (LLMs). These models have performed state-of-the-art downstream tasks such as complex reasoning and QA[6]. Research indicates that models with larger parameters effectively serve as implicit knowledge bases, offering superior integration and application of knowledge compared to traditional external knowledge bases and exhibiting emergent abilities such as in-context learning and instruction following[7, 8]. However, some studies observed limitations in applying LLMs to traditional NLP tasks[9, 10, 11]. For instance, in the named entity recognition task, the performance still lags significantly behind the current best NER model[9]. Additionally, effectively handling datasets for STC tasks, such as Ohsumed and TagMyNews, continues to be challenging. Therefore, our study aimed to leverage the inherent knowledge and emergent capabilities of large models to address this challenge.
A notable approach to enhancing the utilization of knowledge and emergent abilities in LLMs is Chain-of-Thought (CoT) prompting[12]. This technique effectively prompts LLMs to exhibit reasoning processes by necessitating the articulation of intermediate problem-solving steps. Consequently, our study focuses on fostering CoT capabilities within LLMs to facilitate the generation of rationales that directly navigate the intricacies inherent in STC tasks. Furthermore, we seek to foster similar capabilities in smaller language models by transferring knowledge from LLMs, particularly in contexts with limited computational resources or when costly API calls are required to access CoT-capable LLMs.
@@
Specifically, we propose Quartet Logic: A Four-Step Reasoning (QLFR) framework. This framework mainly includes Syntactic and Semantic Enrichment CoT (SSE-CoT), which decomposes the STC task into four more manageable subtasks: (i) key-concept identification, which identifies critical words in the input text; (ii) common-sense knowledge retrieval, which retrieves common-sense knowledge pertinent to these keywords, bridging semantic gaps in short texts; (iii) text rewriting, which rephrases the texts using this knowledge to enhance syntax and readability; and (iv) short-text classification, which employs these refined texts for classification. This strategic design effectively addresses the STC challenges through guided CoTs.
@@
Additionally, we expanded our framework to propose the method CoT-Driven Multi-task learning (QLFR-CML) method, which boosts smaller models by transferring task-aware knowledge from LLMs. This approach aligns closely with practical application requirements for real-world scenarios. In this approach, we extract task-specific rationales from both SSE-CoT, which acquires common-sense knowledge, and Domain Augmentation CoT (DA-CoT), which garners domain knowledge. Subsequently, we employ multi-task learning to fine-tune the smaller models by integrating the ground truth with the rationales from these two distinct CoTs. Moreover, our Explicit Category Context Augmentation (ECCA) strategy refines the model’s output by aligning it more closely with the ground truth.
@@
    To the best of our knowledge, this research is the first to employ Large Language Models alongside chain-of-thought reasoning to investigate and tackle challenges in short-text classification STC tasks.
@@
    We introduce Quartet Logic: A Four-Step Reasoning (QLFR) framework incorporating a novel Syntactic and Semantic Enrichment Chain-of-Thought (SSE-CoT) approach. This approach enables LLMs to effectively decompose STC tasks and address semantic sparsity and syntactic ambiguity by breaking them into more manageable subtasks.
@@
    We propose a CoT-Driven Multi-task learning (QLFR-CML) method designed to enhance the performance of smaller models for STC tasks. This framework leverages task-aware knowledge transfer from LLMs to address the inherent challenges in STC tasks, which can alleviate the challenges of STC tasks in real-world scenarios.
@@
    Comprehensive experiments were conducted using six challenging short-text benchmark datasets. The experiments confirm that our QLFR framework can significantly utilize LLMs for this challenging task and is superior to several novel baselines. QLFR-CML also shows its capacity to enhance smaller models, even with limited computational resources.
@@ 2.1 Text Classification
Text classification (TC) is a pivotal task in Natural Language Processing NLP that assigns predefined labels to text entities[13]. Historically, TC has been performed through a two-stage process: feature extraction using techniques such as term frequency-inverse document frequency[14], followed by applying classifiers such as support vector machines[15]. Deep learning has revolutionized this approach, allowing for an end-to-end methodology. Contemporary models, including convolutional neural networks[16] and long short-term memory networks[17], can learn independently from raw data, bypass manual feature engineering, and enhance classification adaptability. The introduction of BERT[18] marked a new phase in deep learning and has become a prevalent choice for TC research.
Graph neural networks (GNNs) have emerged as a significant development in TC. Some GNNs, such as TLGNN[19], TextING[20], and HyperGAT[21], treat each document as a network of interconnected word nodes, effectively transforming the text classification challenge into a graph-based task. Although this method introduces a novel viewpoint, its effectiveness can diminish in situations with limited labeled data. On the other hand, models like TextGCN[2] and TensorGCN[22] take a broader approach, framing the classification task within corpus-level graphs where both individual words and entire texts serve as nodes. These models use node classification techniques to identify and classify unlabeled textual elements. However, they may falter when handling concise textual data or data with limited contextual richness.
Recently, the introduction of ChatGPT has revolutionized the field. Numerous studies have used large Language Models for TC tasks. Some studies have leveraged ChatGPT for automated genre recognition to streamline the text classification process through its zero-shot classification abilities[23] and assessed the capacity of ChatGPT for text classification within affective computing by employing it for personality prediction, sentiment analysis, and suicidal ideation detection tasks[24].
@@ 2.2 Short Text Classification
Short Text Classification (STC) has consistently emerged as a significant obstacle in the field of NLP, presenting unique and complex challenges in sharp contrast to those encountered in traditional TC tasks[25, 13, 26]. First, the brevity of a text inherently limits its semantic and syntactic integrity[1, 5]. Scholars have sought methods to augment the expressiveness of short texts by integrating additional information[27, 28, 29]. Commonly employed instances include concepts found in external knowledge bases such as [30] and latent topics uncovered within the corpus[31, 32]. Second, STC frequently encounters the dilemma of sparsely labeled data in practical applications[4, 33], thereby increasing the task complexity. A prevailing strategy to mitigate this entails employing graph-based methods, which not only supplement additional information but also offset the paucity of label data, as evidenced by the approach adopted by [34], constructing a corpus-level graph that models latent topics, entities, and documents jointly, where the entities are words linked to knowledge graphs. SHINE[35], which introduces a hierarchically organized heterogeneous corpus-level graph consisting of word-level and document-level graphs, can fully exploit interactions between nodes of the same types and capture similarities between short texts. ST-Text-GCN[36] uses a self-training method for keyword extraction to fully utilize limited labeled texts and a large number of unlabeled texts.
@@
Despite the advancements in graph-based methodologies, they exhibit clear limitations in specific practical scenarios. One of the primary drawbacks is the necessity to retrain the entire model to incorporate new test samples, which can be both computationally intensive and time-consuming. Consequently, there has been a growing interest in inductive reasoning approaches, which obviate the need for retraining by integrating new samples directly with existing training and unlabeled data. Innovations, such as HGAT-inductive[37], propose a novel framework for inductively connecting each new sample to an existing corpus, thereby facilitating dynamic learning. SimpleSTC[38] adopts a word-only approach to navigate the inductive STC problem accurately. Nevertheless, although these inductive methods offer increased flexibility and efficiency, they may necessitate a trade-off in terms of precision because they rely on the assumption that new samples share similar characteristics with the existing corpus.
2.3 Large Language Models
Large Language Models (LLMs), including the ChatGPT[39], LLama2[40], and ChatGLM series[41], have attracted significant attention due to their innovative dialogue system breakthroughs and their potential for various downstream applications. These LLMs have excelled in embedding extensive common knowledge and demonstrating diverse capabilities[7, 8]. The CoT, which enhances LLMs’ reasoning capacities of LLMs by integrating a systematic, stepwise reasoning process within demonstrations, has been applied to various complex reasoning tasks[12, 42, 43].
Presently, research focusing on LLM over the CoT framework principally navigates through two major trajectories. Initially, emphasis was placed on devising sophisticated, human-like cognitive design strategies with methods such as the Tree of Thought[44], which seamlessly linearizes thought processes to facilitate the retrospective exploration and probing of various inferential branches emanating from a central idea. In turn, the Graph of Thought[45] evolves the tree structure into a directed acyclic graph, embedding self-circulation. In contrast, the array of thoughts[46] maintains a single, perpetually evolving context chain, thereby obviating the need for redundant inquiries within the cognitive tree. Subsequently, investigations were conducted to determine the optimal strategies for enhancing the performance of the LLM in assorted tasks. For example, [47] progressively extracted aspects and opinions to navigate the complexities of implicit sentiment analysis, [10] enhanced the relation extraction with heightened precision.
In recent years, the burgeoning scale of deep learning models has been remarkable, with architectures such as transformers and MOE-facilitating models that exceed a trillion parameters, particularly within the era of Large Language Models (LLMs). While beneficial for performance, this escalation in model size has significantly increased deployment costs and impacted inference efficiency, elevating knowledge distillation to a critical role in model compression. Knowledge distillation[48], which is a technique designed to transfer insights from a voluminous teacher model to a more compact student model, has traditionally focused on imparting general knowledge to LLMs and their smaller counterparts, a process exemplified by approaches such as MINILLM[49] and GKD[50]. Moving beyond conventional methods, some strategies leverage the emergent capabilities of LLMs, such as GPT-3[39], which exhibits unique abilities that can be harnessed through in-context learning, chains of thought, and explicit instruction following. Techniques such as the CoT method incorporate intermediate reasoning into prompts, enhancing models under small-sample learning paradigms, such as Meta-Context Tuning and Multitask-Context Tuning, as seen with MT-COT[51]. Furthermore, approaches such as the IF method, as demonstrated by Lion[52], exploit LLMs’ adaptability of LLMs to generate challenging instructions that fortify the capabilities of student models. These innovative methods provide new insights into knowledge distillation and expand the horizons for model compression, marking a significant stride in the efficient utilization of LLMs.

### v2

@@ Abstract
Short Text Classification (STC) is crucial for processing and understanding the brief but substantial content prevalent on contemporary digital platforms. The STC encounters difficulties in grasping the semantic and syntactic intricacies, an issue that is apparent in traditional pre-trained language models. Although Graph Convolutional Networks enhance performance by integrating external knowledge bases, these methods are limited by the quality and extent of the knowledge applied. Recently, the emergence of Large Language Models (LLMs) and Chain-of-Thought (CoT) has significantly improved the performance of complex reasoning tasks. However, some studies have highlighted the limitations of their application in fundamental NLP tasks. Consequently, this study first employs CoT to investigate and enhance the capabilities of LLMs in STC tasks. We propose the Syntactic and Semantic Enrichment CoT (SSE-CoT) method, effectively decomposing the STC tasks into four distinct steps: (i) essential concept identification, (ii) common-sense knowledge retrieval, (iii) text rewriting, and (iv) classification. Furthermore, recognizing resource constraints in sectors like finance and healthcare, we then introduce the CoT-Driven Multi-Task Learning (CDMT) framework to extend these capabilities to smaller models. This framework begins by extracting rationales from LLMs and subsequently fine-tunes smaller models to optimize their performance. Extensive experimentation across six short-text benchmarks validated the efficacy of the proposed methods. In particular, SSE-CoT achieved state-of-the-art performance with substantial improvements on all datasets, particularly on the Ohsumed and TagMyNews datasets.
@@ 1 Introduction
Short texts are crucial to the contemporary flow of information, particularly with the rapid growth of the Internet[1]. They play an essential role on major social media platforms, including Twitter, TikTok, Instagram, and Weibo, where they facilitate social interaction and are integral to daily activities. As a critical task for intelligent empowerment and the application of short texts, short text classification (STC) is essential for applications such as news categorization[2], question answering (QA)[3], and sentiment analysis[4]. Traditional pre-trained language models (PLMs) struggle with the semantic and syntactic intricacies of STC[1, 5]. As shown in Fig 1, categorizing a news title such as ‘Del Potro says make French Open’ into the ‘sport’ category can be challenging, particularly without recognizing ‘Del Potro’ as a professional tennis player. The absence of a clear subject and predicate further hinders the model’s understanding and learning capacity. To address the challenges inherent in STC tasks, Graph Convolutional Networks (GCNs) have shown some progress by incorporating an additional knowledge base and redefining STC as a node classification issue, which mitigates the problem of limited training data. However, the effectiveness of GCNs is remains constrained by the quality and scope of the knowledge employed.
@@
Figure 1: This diagram compares two approaches to the STC tasks. Due to misinterpretation, the traditional approach erroneously classifies the input ‘Del Potro says make French Open’ as ‘world’. Conversely, our CoT method employs a sequential analytical process that correctly identifies ‘Del Potro’ as a tennis player, recognizes the ‘French Open’ as a tennis tournament, and detects the absence of a grammatical object in the sentence, resulting in accurate categorization under ‘sport’.
@@
The NLP landscape has been revolutionized by Large Language Models (LLMs), which have achieved state-of-the-art performance in downstream tasks such as complex reasoning and QA[6]. Research suggests that models with larger parameters effectively function as implicit knowledge bases, offering superior integration and application of knowledge compared to traditional external knowledge bases, while demonstrating emergent abilities such as in-context learning and instruction following[7, 8]. However, some studies have highlighted limitations in applying LLMs to traditional NLP tasks[9, 10, 11]. For example, in the task of named entity recognition, the performance remains significantly below that of the current best NER model[9]. However, the application of LLMs to STC tasks remains unexplored, a gap this study aims to address. Additionally, effectively handling datasets for STC tasks, such as Ohsumed and TagMyNews, remains to be challenging.
Our study begins by examining the challenges faced by LLMs in STC tasks. Typically, these models are pre-trained on extensive text corpora that often fail to effectively capture the semantic and syntactic nuances of short texts, thereby reducing their effectiveness. To address this issue, we utilize Chain-of-Thought (CoT) prompting[12] to improve LLMs performance in STC tasks by enabling step-by-step reasoning. Futhermore, practical challenges arises with LLMs, particularly in fileds such as finance and healthcare, where computational resources are insufficient for fine-tuning domain-specific LLMs. Smaller available models, which lack sufficient knowledge and CoT capabilities, fail to achieve comparable performance to LLMs. Accordingly, the second part of our research focuses on transferring knowledge and CoT capabilities from LLMs to smaller models, enabling them to perform effectively in resource-constrained environments.
@@
Specifically, we propose the Semantic and Syntactic Enrichment CoT (SSE-CoT) method, designed to enhance the performance of LLMs in STC tasks. The framework divides STC tasks into four subtasks: (i) key-concept identification, which involves identifying critical words in the input text; (ii) common-sense knowledge retrieval, facilitating the acquisition of common-sense knowledge relevant to these identified keywords, thereby bridging semantic gaps in short texts; (iii) text rewriting, which reformulates the texts using this acquired knowledge to improve syntax and readability; and (iv) short-text classification, which leverages the refined texts for accurate classification.
@@
Additionally, we introduce the CoT-Driven Multi-Task learning (CDMT) framework, aimed at enhancing smaller models [^1] In our study, we define models with billions of parameters as large models, while those with millions of parameters are classified smaller models. by transferring knowledge and CoT abilities from LLMs. In this framework, we first extract task-specific rationales from both SSE-CoT and Domain Augmentation CoT (DA-CoT). Subsequently, we employ multi-task learning to fine-tune the smaller models using three distinct supervision signals: the rationales from SSE-CoT and DA-CoT, as well as the ground truth. Furthermore, our Explicit Category Context Augmentation (ECCA) strategy enhances model performance by aligning predictions more closely with the ground truth.
@@
    To the best of our knowledge, this research is the first to employ Large Language Models alongside chain-of-thought reasoning to investigate and tackle challenges in short text classification tasks.
@@
    We propose the Syntactic and Semantic Enrichment CoT (SSE-CoT) method to enhance the performance of LLMs for STC tasks. This approach enables LLMs to effectively decompose STC tasks and address semantic sparsity and syntactic ambiguity by breaking them into four subtasks.
@@
    We introduce the CoT-Driven Multi-Task learning (CDMT) framework to improve the capabilities of smaller models in STC tasks. This framework transfers knowledge and CoT abilites from LLMs to smaller models to boost their performance.
@@
    Comprehensive experiments were conducted using six challenging short-text benchmark datasets. The experiments confirm that our SSE-CoT method can significantly utilize LLMs for this challenging task and is superior to several novel baselines. CDMT also shows its capacity to enhance smaller models, even with limited computational resources.
@@ 2.1 Text Classification
Text classification (TC) is a fundamental task in Natural Language Processing NLP that involves assigning predefined labels to text entities[13]. Historically, TC has been approached as a two-stage process: feature extraction using techniques such as term frequency-inverse document frequency[14], followed by the application of classifiers such as support vector machines[15]. The advent of deep learning has revolutionized this approach, enabling an end-to-end methodology. Contemporary models, such as convolutional neural networks[16] and long short-term memory networks[17], learn directly from raw data, bypass manual feature engineering, and improve classification adaptability. The introduction of BERT[18] marked a significant milestone in deep learning, establishing itself as a prevalent choice for TC research.
Graph neural networks (GNNs) have emerged as a pivotal development in TC. Certain GNNs, such as TLGNN[19], TextING[20], and HyperGAT[21], represent each document as a network of interconnected word nodes, effectively reframing the text classification challenge into a graph-based task. While this approach offers a novel perspective, its effectiveness diminish when dealing with limited labeled data. Conversely, models such as TextGCN[2] and TensorGCN[22] adopt a broader perspective, framing the classification task within corpus-level graphs, where both individual words and entire texts are presented as nodes. These models use node classification techniques to identify and classify unlabeled textual elements. However, these models often struggle when handling concise textual data or datasets with limited contextual richness.
Recently, the introduction of ChatGPT has revolutionized the field. Numerous studies have explored the application of Large Language Models in TC tasks. Some studies have leveraged ChatGPT for automated genre recognition to streamline the text classification process through its zero-shot classification abilities[23] and evaluated the capacity of ChatGPT for text classification within affective computing by employing it in tasks such as personality prediction, sentiment analysis, and suicidal ideation detection tasks[24].
@@ 2.2 Short Text Classification
Short Text Classification (STC) has consistently posed significant challenges in the field of NLP, presenting unique complexities that sharply contrast with those of traditional TC tasks[25, 13, 26]. First, the brevity of short texts inherently limits its semantic and syntactic richness[1, 5]. Researchers have explored methods to enhance the expressiveness of short texts by integrating additional information[27, 28, 29]. Common approaches include concepts from external knowledge bases such as [30] and latent topics uncovered within the corpus[31, 32]. Second, STC often encounters the challenge of sparsely labeled data in practical applications[4, 33], which exacerbates increasing the task complexity. A prevalent strategy to mitigate this involves employing graph-based methods, which not only supplement additional information but also offset the paucity of label data, as evidenced by the approach adopted by [34], constructing a corpus-level graph that models latent topics, entities, and documents jointly, where the entities are words linked to knowledge graphs. SHINE[35] introduces a hierarchically organized heterogeneous corpus-level graph, comprising word-level and document-level graphs, to fully exploit interactions between nodes of the same type and capture similarities between short texts. ST-Text-GCN[36] uses a self-training method for keyword extraction, effectively leveraging limited labeled texts and a large number of unlabeled texts.
@@
Despite significant advancements in graph-based methodologies, these approaches exhibit clear limitations in certain practical scenarios. One of the primary drawbacks is the need to retrain the entire model to incorporate new test samples, which can be both computationally intensive and time-consuming. As a result, there has been a growing interest in inductive reasoning approaches, which eliminate the need for retraining by integrating new samples directly with existing training and unlabeled data. Innovations such as HGAT-inductive[37], propose a novel framework for inductively linking each new sample to an existing corpus, thereby facilitating dynamic learning. SimpleSTC[38] adopts a word-only approach to address the inductive STC problem accurately. Nevertheless, while these inductive methods offer increased flexibility and efficiency, they may necessitate a trade-off in terms of precision as they rely on the assumption that new samples share similar characteristics with the existing corpus.
2.3 Chain-of-thought in LLMs
Large Language Models (LLMs), such as [39, 40, 41] have gained significant attention for their advancements in dialogue systems and potential across various applications[7, 8]. The Chain of Thought (CoT) strategy improves the reasoning capabilities of LLMs by employing a structured, step-by-step process suitable for complex reasoning tasks[12, 42, 43].
Recent studies have explored the use of the CoT approach to further enhance the capabilities of LLMs. For instance, [12] improved learning and reasoning in LLMs by manually constructing CoT prompts with specific examples to facilitate the analysis of complex problems. [44] incorporated programming languages as annotated rationales in their PAL method, converting problem-solving into executable Python programs and demonstrating CoT’s utility in programmed tasks. Additionally, [45] employed the CoT strategy with a three-step prompting principle, effectively inferring the latent intent of opinions to address Implicit Sentiment Analysis. [10] applied the CoT strategy to fine-tuning GPT-3 and Flan-T5 enhancing the processing of complex semantic relationships in Relation Extraction tasks. [46] introduced GeM-CoT, a generalizable CoT mechanism designed to enhance performance and generalization across diverse mixed-task scenarios.
Various CoT modifications have been proposed to optimize reasoning processes. The Tree of Thought [47] linearizes reasoning for retrospective analysis, while the Graph of Thought [48] restructures it into a directed acyclic graph textcolorredto enhance navigation. The Array of Thoughts [49] maintains a dynamic context chain to minimize repetitive querying. Although LLMs have been explored for certain NLP basic tasks[9, 10, 11], their application in STC tasks using Chain-of-Thought strategy remains unexplored until our study, motivating the focus of our study. Therefore, this paper explores the integration of CoT with LLMs for addressing STC tasks.
2.4 Knowledge Distillation in LLMs
In recent years, LLMs have undergone significant evolution, with their parameter sizes have expanded significantly. Research suggests that models with larger parameters usually achieve better performance[39]. However, the increasing scale of these models presents challenges, such as higher deployment costs and decreased training efficiency. Knowledge distillation [50] is a method that compresses models by transferring knowledge from a large, well-trained teacher model to a smaller student model.
Traditional knowledge distillation methods using LLMs, such as GKD[51] and MINILLM[52], primarily focus on distilling specific outputs from the teacher model, like classification labels. These methods directly transfer direct knowledge from the teacher to the student model, simplifying complex concepts to facilitate easier learning for the smaller model. In contrast, the approach presented in this paper emphasizes distilling the teacher model’s thought process in response to inputs, capturing the reasoning underpinning its decisions. Recent advancements, notably MT-COT[53], integrate nuanced elements of the teacher’s thought process into the student model’s training. This enhances the student’s ability to manage complex tasks and multitask through strategically crafted prompts. Furthermore, SOCRATIC COT [54] distills reasoning capabilities from the teacher model into the student model to solve complex problems.